### v1

# Correctness Assessment of Code Generated by Large Language Models Using Internal Representations

Abstract
Ensuring the correctness of code generated by Large Language Models (LLMs) presents a significant challenge in AI-driven software development. Existing approaches predominantly rely on black-box (closed-box) approaches that evaluate correctness post-generation, failing to utilize the rich insights embedded in the LLMs’ internal states during code generation. In this paper, we introduce Openia, a novel white-box (open-box) framework that leverages these internal representations to assess the correctness of LLM-generated code. Openia systematically analyzes the intermediate states of representative open-source LLMs specialized for code, including DeepSeek-Coder, CodeLlama, and MagicCoder, across diverse code generation benchmarks. Our empirical analysis reveals that these internal representations encode latent information, which strongly correlates with the correctness of the generated code. Building on these insights, Openia uses a white-box/open-box approach to make informed predictions about code correctness, offering significant advantages in adaptability and robustness over traditional classification-based methods and zero-shot approaches. Experimental results demonstrate that Openia consistently outperforms baseline models, achieving higher accuracy, precision, recall, and F1-Scores with up to a 2X improvement in standalone code generation and a 46% enhancement in repository-specific scenarios. By unlocking the potential of in-process signals, Openia paves the way for more proactive and efficient quality assurance mechanisms in LLM-assisted code generation.
Index Terms:
LLM-generated code, code LLMs, code quality assessment, internal representations, white-box, open-box approach
1 Introduction
The emergence of Large Language Models for Code (Code LLMs) has transformed the landscape of automated code generation, offering a promising solution to improve developer productivity and address the growing demand for software [1, 2, 3, 4]. These models, trained on vast repositories of code, are capable of generating complex functionalities, suggesting optimizations, and even automating repetitive tasks. However, as LLM-generated code becomes increasingly integrated into real-world systems, ensuring the quality and reliability of this code has become critical [5, 6, 7, 8]. Buggy/vulnerable code can lead to significant functional failures, increased maintenance costs, and severe security vulnerabilities, emphasizing the need for rigorous evaluation and assurance mechanisms [9].
Despite this, recent studies show a concerning trend that LLM-generated code frequently contains more bugs and security vulnerabilities than human-written code [10, 11, 12, 7, 13, 8, 14, 5]. This problem has led developers to find LLM-generated code often unusable, resulting in considerable time spent debugging or discarding it entirely [15, 1]. Current approaches to detect and mitigate these errors rely heavily on post-hoc analyses, such as static or dynamic code analysis and testing, which can be resource-intensive and may miss context-specific issues.
Human coding processes remain a “closed-box”, where we only see the final code, with no insights into the developer’s thought process or intermediate steps. In contrast, AI (LLMs) code generation is inherently accessible: we can examine the model’s internal states and byproducts, especially with open-source code LLMs [16, 17, 18], providing a transparent, “open-box” advantage unique to code LLMs. This visibility opens an exciting opportunity: Can we leverage this visibility to proactively detect and correct issues during the code generation process, preventing bugs and vulnerabilities before they reach developers?
While extensive research has been conducted on the performance and capabilities of code LLMs [8, 19, 20, 21, 22, 7], there is a notable gap in understanding how their internal representations, specifically the hidden states during code generation, encode information related to reliability. These internal states capture the model’s intermediate understanding of the input and output, but their relationship with code quality remains largely unexplored. If these representations do indeed encode reliability signals, they could offer a novel pathway to assess and enhance code quality directly during generation rather than relying solely on external validation.
We hypothesize that the internal states of LLMs encapsulate meaningful signals that reflect the reliability of the code they generate. These signals, embedded in the representations formed during the generation process, may correlate with key quality attributes. Understanding and leveraging these signals could enable the proactive identification of faults, thereby improving the overall quality assurance pipeline.
In this work, we tackle the challenges and gaps in assessing the reliability of code generated by LLMs by presenting two key contributions. First, we introduce Openia, a novel framework that leverages the internal representations of code LLMs during code generation to assess the correctness of the generated code. Unlike traditional quality assurance techniques that rely heavily on post-hoc validation, Openia identifies and utilizes the signals embedded within the LLM’s internal states. By doing so, it enables earlier detection of potential bugs, reducing both the computational cost and the time required for quality assurance. This proactive approach enhances the utility of LLMs in generating high-quality code for real-world applications. Second, we conduct a systematic analysis of the internal signals of code LLMs during code generation, focusing on three representative open-source models, including DeepSeek Coder, Code Llama, and Magic Coder. These models are evaluated across diverse benchmarks, encompassing standalone code generation and repository-level generation tasks. Our analysis reveals specific patterns and features in the internal representations that align with the reliability attributes of the generated code. Specifically, our experimental results demonstrate that Openia consistently outperforms traditional classification-based and zero-shot approaches across a range of benchmarks. For independent/standalone code generation tasks, Openia achieves up to a 2X improvement in accuracy compared to zero-shot baselines and surpasses classification-based methods, such as those utilizing CodeBERT and CodeT5+, by up to 20% in accuracy. In repository-level code generation, Openia shows a remarkable, up to 46%, improvement in F1-Score, highlighting its robustness in handling complex and context-dependent coding scenarios. These findings provide valuable insights into how LLMs encode and process information related to code correctness, offering a deeper understanding of their inner workings.
By combining these contributions, this work lays the foundation for more efficient and proactive approaches to ensuring the quality of LLM-generated code. It highlights the latent capabilities of code LLMs to encode reliability-related information and demonstrates the potential of leveraging these signals to enhance quality assurance in automated software development.
2 Internal Representations in LLMs for Code Generation
Code LLMs rely on a sequence-to-sequence mapping mechanism, where input sequences (e.g., natural language descriptions or partially written code) are mapped to output sequences (e.g., complete or corrected code) [23, 19, 20, 24]. Central to this process is the internal representations, which encode the intermediate understanding of the model as it processes inputs and generates outputs. These representations provide a window into the LLM’s decision-making process, offering insights into how the model predicts the next code tokens.
2.1 The Code Generation Process
For a code LLM $\mathscr{M}$, given an input sequence $x=\{x_{1},x_{2},...,x_{n}\}$, where $x_{i}$ represents the $i$-th token in the input, $\mathscr{M}$ generates an code output sequence $c=\{c_{1},c_{2},...,c_{m}\}$ token by token. At each generation step $s$, the model predicts the next code token $c_{s}$ by leveraging both the input sequence and the tokens generated so far, denoted $c_{<s}=\{c_{1},c_{2},...,c_{s-1}\}$. The prediction process at step $s$ can be represented as:
$P(c_{s}|x,c_{<s})=\text{softmax}(W_{o}\cdot h_{L,s}+b_{o})$
where $h_{L,s}$ is the final hidden state of the LLM at layer $L$ for token $c_{s}$, $W_{o}$ and $b_{o}$ are the output layer weights and bias.
2.2 Internal Representations
Internal representations, or intermediate activations, are vectors $h_{l,s}$ that capture the latent information processed by the model at a specific layer $l$ for a given code token $c_{s}$. These activations are computed as:
$h_{l,s}=f_{l}(h_{l-1,s},\text{context}_{s})$
where $h_{l-1,s}$ is the activation from the preceding layer $(l-1)$. Meanwhile, $\text{context}_{s}$ represents contextual information (e.g., attention-weighted representations of other tokens), $f_{l}$ is the layer-specific transformation function (e.g., feedforward, attention, or normalization operations) applied at layer $l$. Note that the initial hidden state for token $c_{s}$ is $h_{0,s}$ is the embedding vector $e_{s}$ of $c_{s}$.
Each $h_{l,s}$ is a vector $\mathbb{R}^{d}$ where $d$ is the dimensionality of the LLM’s hidden states (e.g., 4,096 in the case of Code Llama). The sequence of activations across all layers and code tokens forms a hierarchical representation of the input and generated sequences, that could encode syntactic, semantic, and contextual information at varying levels of abstraction. Internal representations encapsulate the underlying structure and logic the model uses to generate code. By analyzing these representations, we gain insights into the model’s ”thought process“ and ”intuition” during code generation. These representations offer valuable cues about the correctness and reliability of the generated code $c$.
3 A Framework Leveraging Internal Representations for LLM Generated Code Quality Assessment
3.1 Task Formulation
Given a code LLM $\mathscr{M}$, an input prompt $x$, and the corresponding LLM-generated code $c$, the objective is to predict whether $c$ meets the correctness criteria as defined by the requirements of $x$. The correctness of $c$ is evaluated by verifying that it aligns with the intended behavior and satisfies the specified requirements provided in $x$.
Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represent a set of validation tests designed to assess whether $c$ adheres to the requirements in $x$. Each test $t\in\mathcal{T}$ returns an outcome, denoted as $t(c)$, which indicates whether $c$ passed or failed the test. Code $c$ is considered correct if and only if it passes all the tests in $\mathcal{T}$.
$\text{Correctness}(c)=\begin{cases}1&\text{if }\forall t\in\mathcal{T},\;t(c)=\textit{passed},\\ 0&\text{otherwise}.\end{cases}$
The goal is to develop a predictive model $\mathscr{Q}$ that evaluates the correctness of generated code $c$ based on the requirements specified by the input prompt $x$ and the context provided by the code LLM $\mathscr{M}$. Formally:
$\mathscr{Q}(\mathscr{M},x,c)=y$
where $y\in\{0,1\}$ is the predicted correctness label for $c$, with $y=1$ indicating that $c$ meets the correctness criteria and $y=0$ otherwise.
3.2 Framework Overview
Figure 1: Openia’s overview
Fig. 1 illustrates the overview of our framework, Openia, which leverages the internal representations of the LLM to assess the correctness of its generated code. Specifically, given a code LLM $\mathscr{M}$ and its output code sequence $c=\{c_{1},c_{2},...,c_{m}\}$, Openia extracts the all internal states produced during $\mathscr{M}$’s generation:
$H=\big{\{}h_{l,s}|l\in\{1,...,L\},s\in\{1,...,m\}\big{\}}$
where $h_{l,s}$ represents the hidden state at layer $l$ and token $c_{s}$. These internal representations are then fed to a probing classifier [25] that predicts whether the generated code $c$ is correct or secure.
Layer Selection for Internal Representations. Modern LLMs often contain multiple transformer layers, each contributing uniquely to the model’s ability to understand and process various tasks. For example, DeepSeek-Coder-6.7B [16] and Code Llama-7B [17] comprise 32 layers, while the 34B version of Code Llama has 60 layers. Each layer captures different levels of abstraction and semantic representation; shallow layers often encode lower-level syntactic and lexical features, while deeper layers focus more on abstract and semantic representations. Jin et al.  [26] have demonstrated the alignment of shallow and deep layers with simpler and more complex tasks, respectively. Meanwhile, Song et al.  [27] highlights the varying importance of layers within an LLM. Thus, selecting the appropriate layer(s) among $L$ layers of $\mathscr{M}$ for extracting internal representations is crucial for achieving optimal performance in downstream tasks. An inappropriate choice of layer may fail to capture the necessary information, leading to suboptimal results. In this work, we conduct several experiments to systematically evaluate the impact of layer selection on the task of assessing the quality of the LLM’s generated code (Sec. 5.2.1).
Token Selection for Internal Representations. In addition to layer selection, the choice of specific tokens among the input tokens in $x$ or generated code tokens in $c$ to extract the internal representations is also very important. Several works [28, 29, 30] have highlighted the effectiveness of probing the final token of the generated answer to assess the answer’s correctness. This token typically encapsulates the accumulated context and semantics of the entire generated sequence. Therefore, the hidden states of this token can provide valuable insights for determining whether the answer is correct or not. Meanwhile, some other approaches [31, 32] focus on the final token of the prompt, just before the response is generated. The reasoning behind this choice is that this token often reflects the model’s transition from input comprehension to output generation, and its hidden states can provide valuable signals into how well the model has understood the task requirement and its readiness to generate the output. In this work, we systematically investigate the impact of different token selection strategies on assessing the correctness of the generated code. Our goal is to understand how token-level choices affect the effectiveness of Openia and to explore the relationship between token-specific hidden states and the correctness of the generated code (Sec. 5.2.2).
4 Evaluation Design
In this research, we seek to answer the following research questions:
  • RQ1. Performance Analysis: To what extent, can Openia with the internal representations of the LLM help identify the correctness of its generated code? How accurately does Openia compare to the baseline approaches?
  • RQ2. Layer and Token Selection Analysis: How do the representations at various token positions and layer depths impact Openia’s performance? Is there any correlation between the complexity of the tasks and the impact of layer and token selection?
  • RQ3. Sensitivity Analysis: How do different factors such as programming languages and task difficulty levels impact Openia’s performance?
  • RQ4. Time Complexity: What is Openia’s running time?
4.1 Dataset Construction
Openia constructs the dataset using four popular benchmarks, HumanEval [33], MBPP [34], and DevEval [35]. In particular, HumanEval and MBPP contain 164 and 500 stand-alone programming problems, respectively, while DevEval consists of 1,825 tasks of repository-level code generation. For each task in these benchmarks, we employed each studied model to generate 10 candidate solutions. We then evaluate the correctness of the generated solutions. To access correctness, we utilized the test cases provided by the corresponding benchmarks. Table I summarizes the number of functional correct and incorrect solutions generated by each model for benchmarks. The full set of the generated code, the corresponding internal representations from the studied code LLMs, and labels could be found on our website [36].
TABLE I: The numbers of correct (#Cor.) and incorrect (#Inc.) solutions of benchmarks
Code LLM HumanEval MBPP DevEval
DeepSeek-1.3B #Cor. 872 2,065 2,092
#Inc. 768 2,701 15,237
DeepSeek-6.7B #Cor. 1,294 3,033 4,321
#Inc. 346 1,749 13,009
CodeLlama-7B #Cor. 644 1,942 3,467
#Inc. 996 2,686 13,863
MagicCoder-7B #Cor. 1,198 3,135 4,113
#Inc. 440 1,752 13,217
4.2 Studied Code LLMs
In this work, experiments are conducted on three representative open-source LLMs specialized for code, including DeepSeek-Coder [16], CodeLlama [17], and MagicCoder [18]. These models were chosen due to their widespread adoption and strong performance in various code-related tasks, making them ideal candidates for analyzing how internal LLM signals correlate with the reliability of generated code. Each model represents a state-of-the-art approach, utilizing architectures and training strategies specifically designed to handle the complexities of programming languages, such as syntax, semantics, and logical consistency. To accommodate hardware resource limitations, we focus on versions of these models with parameter sizes not exceeding 7B. Despite their smaller sizes, these models retain robust capabilities for code generation, ensuring that the results of our study remain relevant to practical applications. This selection also enables a fair comparison by standardizing computational resources across experiments. For consistency and reproducibility, Openia employs their base models with pre-trained weights, which are officially published on HuggingFace. Importantly, we do not perform any fine-tuning on the models to ensure the evaluation reflects their out-of-the-box capabilities.
4.3 Experimental Procedure
RQ1. Performance analysis:
Baselines: We evaluate the performance of Openia by comparing it with two baselines:
  • Classification-based [37, 38, 39]: These approaches focus on capturing the semantics of code to predict its correctness. Specifically, pre-trained models such as CodeBERT [40] and CodeT5 [41]/CodeT5+ [42] are employed to encode the semantics of the code, which is generated by the LLM. The resulting embeddings are then used to train a classifier which determines whether the given code is correct.
  • Prompting-based [43]: These approaches leverage the reasoning and generalization capabilities of LLMs to access the code correctness without requiring additional fine-tuning or training. Specifically, prompting techniques such as zero-shot and few-shot learning are applied to directly query the LLM for predictions on the correctness of the given code.
For a fair comparison, in all classification-based approaches and Openia, we apply the same neural architecture for the classifier containing the input, output layers, and two hidden layers with 128 and 64 neurons correspondingly. And the training process runs for 50 epochs with a batch size of 32, using a learning rate of 0.001.
Procedure: For comparison with the baselines, the internal state of the last generated token in the last layer of the studied LLM is extracted for each query as a representation to assess the correctness of the corresponding generated output. This choice is motivated by the fact that the internal state from the last layer is closest to the model’s final decision, making it a logical and effective representation for evaluating the correctness of the generated code. Moreover, the last token is generated based on both the prompt input and the entire generated output, allowing its internal state to encapsulate the most comprehensive information processed by the model. Therefore, this state is selected for conducting experiments to answer this research question. While this experiment focuses on leveraging the last token’s internal state for comparisons with baselines, the impacts of other tokens and layers are investigated separately in RQ2.
In this experiment, we compare the performance of Openia with the baselines across multiple settings.
  • Independent-unit code generation: This evaluates the correctness of code generated for standalone programming tasks [33, 34].
    • Cross-benchmark: The internal representations of a given code LLM obtained while generating code for tasks in one benchmark (i.e., MBPP) are used to train the probing classifier. The classifier is then tested on the internal representations obtained while generating code for tasks in a different benchmark (i.e., HumanEval) with the same code LLM.
    • Cross-task: The tasks within a benchmark are split in an 9:1 ratio, where 90% of the tasks are used to extract internal representations for training the probing classifier, and 10% of the remaining tasks are used for testing.
  • Repo(sitory)-level code generation: This evaluates the correctness of code generated for existing projects. Unlike independent-unit code generation, this setting involves understanding the broader context of the codebase [44, 45]. This setting is particularly challenging, as it requires LLMs to maintain contextual consistency and correctness across multiple units of code within a project.
RQ2. Layer and token selection analysis: This experiment investigates the impact of internal states from different layers and token positions on Openia’s performance in assessing the correctness of generated code. To answer this research question, we systematically evaluate the performance of Openia when it is trained and tested using internal states extracted from various layers and token positions across multiple settings.
RQ3: Sensitivity analysis: This experiment examines Openia’s ability to generalize across programming languages. Specifically, we evaluate whether Openia, trained on the internal states from the generation of code in several programming languages, can effectively predict the correctness of code in a completely different language. In addition, we also investigate the performance of Openia with code generated for tasks of different difficulty levels.
4.4 Metrics
The task of accessing the correctness of code generated by LLM can be formulated as a binary classification task. Specifically, given a generated code snippet (along with related information, such as its internal representations), Openia and the baseline approaches aim to predict whether the code is correct or not. To evaluate the performance of the approaches, we adapt widely used classification metrics: Accuracy, Precision, Recall, and F1-Score.
Given the potential class imbalance in the dataset, as highlighted in Table I, where the ratio of correct to incorrect code snippets may vary significantly, it is critical to ensure a fair evaluation across both classes. To address this, we employ weighted metrics, including Weighted Accuracy, Weighted Prediction, Weighted Recall, and Weighted F1-score. These metrics compare the corresponding result for each label and return the average considering the proportion for each label in the test set.
5 Experimental Results
5.1 Performance Analysis
5.1.1 Performance Comparison
TABLE II: Correctness assessment performance for independent-unit code generation
Code LLM Setting Accuracy Precision Recall F1-Score
DeepSeek-1.3B Cross-benchmark Zeroshot 0.55 0.62 0.55 0.50
CodeBERT 0.50 0.52 0.50 0.45
CodeT5+ 0.47 0.52 0.47 0.34
Openia 0.66 0.66 0.66 0.66
Cross-task Zeroshot 0.42 0.42 0.42 0.42
CodeBERT 0.64 0.64 0.64 0.63
CodeT5+ 0.66 0.66 0.66 0.66
Openia 0.73 0.73 0.73 0.73
DeepSeek-6.7B Cross-benchmark Zeroshot 0.32 0.69 0.32 0.26
CodeBERT 0.35 0.62 0.35 0.35
CodeT5+ 0.66 0.65 0.66 0.66
Openia 0.65 0.69 0.65 0.67
Cross-task Zeroshot 0.68 0.78 0.68 0.57
CodeBERT 0.64 0.60 0.64 0.61
CodeT5+ 0.69 0.68 0.69 0.68
Openia 0.79 0.79 0.79 0.79
CodeLlama-7B Cross-benchmark Zeroshot 0.61 0.37 0.61 0.46
CodeBERT 0.58 0.54 0.58 0.54
CodeT5+ 0.63 0.62 0.63 0.62
Openia 0.69 0.68 0.69 0.68
Cross-task Zeroshot 0.50 0.25 0.50 0.33
CodeBERT 0.68 0.68 0.68 0.68
CodeT5+ 0.70 0.70 0.70 0.70
Openia 0.69 0.76 0.75 0.75
MagicCoder-7B Cross-benchmark Zeroshot 0.27 0.07 0.27 0.11
CodeBERT 0.46 0.58 0.46 0.49
CodeT5+ 0.54 0.64 0.54 0.56
Openia 0.56 0.67 0.56 0.58
Cross-task Zeroshot 0.48 0.73 0.48 0.47
CodeBERT 0.75 0.73 0.75 0.73
CodeT5+ 0.68 0.68 0.68 0.68
Openia 0.78 0.78 0.78 0.78
Table II and Table III show the performance of correctness assessment for independent-unit code generation and repo-level code generation for the studied code LLMs, DeepSeekCoder 1.3B and 6.7B, CodeLlama-7B, and MagicCoder-7B. Overall, Openia consistently outperforms all baselines across all experimental settings in all evaluation metrics and for all studied code LLMs, demonstrating its robustness in diverse scenarios.
For independent-unit code generation shown in Table II, in Cross-benchmark, Openia achieves an accuracy of up to 0.69 (for CodeLlama-7B), surpassing the baseline approaches by from 9% to 118% relatively. In Cross-task, Openia obtains accuracies of 0.69 (for CodeLlama-7B) to 0.79 (for DeepSeek-6.7B), exceeding the baseline performance by up to 2 times. This superior performance highlights the effectiveness of Openia by leveraging the internal representations of the LLM in assessing the correctness of generated code.
For repo-level code generation shown in Table III, Openia also consistently surpasses the baseline approaches. For example, Openia obtains the F1-Scores from 0.73 to 0.79, while the corresponding figures of Zeroshot, CodeBERT, and CodeT5+ range from 0.50 to 0.71. Especially for CodeLlama-7B, Openia’s accuracy is 0.81, and the improvements by Openia in F1-Score are from 11% to 46%. These results highlight the effectiveness of leveraging internal representations of the LLM, making Openia more reliable in assessing the correctness of the real-world repo-level code.
In our experiments, Zeroshot, which prompts the LLM to determine whether a given code snippet is correct, exhibits an imbalance in handling positive and negative cases. Specifically, while this approach can achieve competitive Precision, it struggles with Recall, especially for DeepSeek Coder variants. For DeepSeek-6.7B, Zeroshot obtains a Precision of 0.69 and 0.70 in independent-unit and repo-level code generation settings, respectively, which is comparable to the other approaches, yet its Recall is up to about two times lower than the others. The reason is that the LLM tends to be over-confident in the correctness of its generated code, frequently labeling its generated code as “correct” even when it is actually incorrect.
Traditional classification-based approaches, which predict code correctness using the code semantics encoded by pre-trained models such as CodeBERT and CodeT5/CodeT5+, demonstrate a significantly stronger performance compared to Zeroshot. Despite their strengths, however, these baselines are outperformed by Openia, as evidenced by consistently higher accuracy, precision, recall, and F1-Scores across all models in Table II and Table III. For repo-level code generation, Openia achieves an F1-Score of 0.73 with DeepSeek-6.7B, compared to 0.69 for CodeT5+. Especially for CodeLlama-7B, the margin of improvement is even more pronounced, with Openia attaining F1-Scores of 0.79, respectively, compared to 0.71 and 0.69 for the classifier based on CodeT5+ and CodeBERT. These results suggest that relying on pre-trained models to encode generated code, while effective to an extent, is insufficient for capturing the nuanced, dynamic, and process-oriented aspects of code correctness.
Unlike these black-box (closed-box) approaches that rely solely on final outputs, Openia is a white-box (open-box) approach that examines the internal workings of the LLM to make predictions. By analyzing the intermediate states of the LLM during code generation, Openia assesses code correctness based on a more comprehensive set of signals derived from these states . This approach uncovers hidden patterns associated with the generated code, mitigating the overconfidence often seen in the Zeroshot approach and providing a more nuanced and reliable evaluation. Moreover, the internal representations utilized by Openia not only encode the code features as CodeBERT and CodeT5+ but also capture the model’s reasoning process. This reasoning process information reflects not just dataset-specific patterns but also fundamental programming underlying programming principles. Openia can dynamically adapt to varying requirements, programming constructs, and output expectations across different tasks and benchmarks, enhancing its generality and robustness. This adaptability enhances its generality and robustness, enabling it to consistently outperform baseline methods across all experimental settings. Additionally, by leveraging these in-process signals, Openia provides a unique advantage: it can detect potential issues and assess correctness as the code is being generated, enabling real-time quality assurance.
TABLE III: Correctness assessment performance for repo-level code generation
Code LLM Accuracy Precision Recall F1-Score
DeepSeek-1.3B ZeroShot 0.24 0.76 0.24 0.27
CodeBERT 0.86 0.78 0.86 0.81
CodeT5+ 0.86 0.81 0.86 0.83
Openia 0.86 0.82 0.86 0.83
DeepSeek-6.7B ZeroShot 0.48 0.70 0.48 0.50
CodeBERT 0.71 0.67 0.71 0.68
CodeT5+ 0.73 0.69 0.73 0.69
Openia 0.75 0.73 0.75 0.73
MagicCoder-7B ZeroShot 0.74 0.64 0.74 0.65
CodeBERT 0.73 0.68 0.73 0.69
CodeT5+ 0.72 0.67 0.72 0.69
Openia 0.77 0.73 0.77 0.73
CodeLlama-7B ZeroShot 0.51 0.65 0.51 0.54
CodeBERT 0.74 0.68 0.74 0.69
CodeT5+ 0.76 0.71 0.76 0.71
Openia 0.81 0.79 0.81 0.79
5.1.2 Code length analysis
Figure 2: Impact of the length of the generated code on Openia’s performance
Fig. 2 illustrates that Openia achieves its best performance with short generated code. Specifically, Openia accurately predicts the correctness of the code snippets with lengths ranging from 1 to 30 tokens, achieving an F1-Score of 0.81. This is because short code snippets are often simple and contain few dependencies. Thus, the internal states of the LLM the semantic and syntactic features of the code can be fully encoded in the internal states of the LLMs, thereby facilitating accurate predictions.
However, the F1-score of Openia decreases as the code length increases, reaching its lowest point (0.66) for medium-length code containing 51-70 tokens. Longer code tends to be more complex, making it challenging for the internal states to fully capture all critical information. This limitation results in a performance drop for Openia.
Interestingly, Openia exhibits performance recovery for longer code, accurately predicting code correctness of exceeding 70 tokens. The reason is that the richer context and redundancy in longer code allow the LLM to better resolve dependencies, reinforce patterns, and effectively focus on relevant sections, leading to improved prediction accuracy. These results highlight the adaptability of Openia in leveraging the LLM’s representations across varying lengths.
5.1.3 Error analysis
TABLE IV: Openia’s performance on different errors
Category F1-Score #Cases
Correct 0.75 644
Timed out 0.89 10
AssertionError 0.82 825
Not defined variables/attributes 0.86 36
Unsupported operands 0.82 13
Index out of range 0.73 19
Type error 0.83 33
Others 0.85 60
Table IV provides insights into Openia’s ability to predict correct and incorrect code by analyzing different error categories. In this table, the Correct row represents Openia’s performance in accurately predicting correct code. Meanwhile, the remaining rows are the results of Openia in detecting incorrect code, categorized by error types.
Overall, Openia demonstrates higher accuracy in detecting incorrect code compared to the correct one. Specifically, Openia’s F1-Score for identifying correct code is 0.75, whereas its performance in detecting incorrect code reaches 0.82. This is reasonable because identifying incorrectness often requires just a single signal/pattern indicating an error, making it easier to classify a code snippet as incorrect. In contrast, determining whether a program is correct requires assessing multiple factors such as logic, structure, and functionality, which makes it a more challenging task.
Among the error types, Openia exhibits the lowest performance in detecting incorrect code caused by index out of range errors. Specifically, its F1-Score for identifying incorrect code of this error is 0.73, which is 12% lower than its average performance of detecting incorrect code. Unlike static errors, such as undefined variables or type errors, that are inherently present in the code and can be directly captured in the LLM’s internal states, index out-of-range errors are not always explicitly encoded in the internal representations of the model. This is because such errors often depend on dynamic factors like runtime conditions or specific input values. Therefore, to effectively detect incorrect code of this error, it could require a deeper understanding of program execution and data dependencies.
5.2 Layer and Token Selection Analysis
5.2.1 Layer selection analysis
(a) Independent-unit code generation
(b) Repo-level code generation
Figure 3: Impact of layer selection on Openia’s performance
Fig. 3 shows the performance of Openia in correctness assessment when using hidden states from different layers of three code LLMs: MagicCoder 7B, CodeLlama 7B, and DeepSeek Coder 6.7B. In general, the representations from the middle layers of the models enable Openia to achieve optimal performance for both independent-unit code generation and repo-level code generation tasks. For independent-unit code generation (Fig. 3(a)), Openia obtains its best performance by leveraging the representations from layer 28 of MagicCoder, layer 16 of CodeLlama, and layer 20 of DeepSeek. Similarly, for repo-level code generation (Fig. 3(b)), Openia attains its highest F1-Score using the representations from layer 28 for both MagicCoder and DeepSeek. In the case of CodeLlama, while Openia’s peak performance is achieved with the representations from the last layer (layer 32), the middle layers (layers 16-28) still offer consistently strong results. These findings suggest that the middle layers in the LLMs typically capture a balance between the local context focused by the shallow layers and the global context encoded by the deep layers, making them optimal for assessing code correctness.
The shallow layers primarily focus on syntax and token-level details, capturing the immediate relationships between adjacent tokens (local context). While such information is useful for understanding the basic structure, it is insufficient for determining the correctness of the generated code. Consequently, leveraging the representations from the shallow layers often leads to low performance of Openia. For example, when using the representations from the first layer of CodeLlama, Openia achieves an F1-Score of 0.59, which is 30% lower than its highest score.
In contrast, the deep layers often capture broader logical flow and long-range dependencies between code tokens (global context). The representations from these layers provide richer semantic information, improving Openia’s ability to assess code correctness. However, the deeper the layers, the closer they are to the model’s final output, making the representations at the deeper layers more task-specific. Leveraging the representations from too deep layers could make Openia to be overfitting with the training data and potentially hinder its generality. For instance, Openia observes a slight decline in performance when using the representations from the last layer of several models.
The middle layers strike a balance between the shallow and the deep layers, providing rich representations that are often suitable for downstream tasks such as code correctness assessment. However, the specific layer at which the performance of Openia peaks or dips can vary depending on the model’s architecture and its pre-training objectives. This suggests that static layer selection may not be ideal, as there is no one-size-fits-all approach to selecting the best layers for Openia. Instead, Openia could benefit from a dynamic approach that adapts layer selection based on the models and task requirements.
5.2.2 Token selection analysis
(a) Independent-unit code generation
(b) Repo-level code generation
Figure 4: Impact of token selection on Openia’s performance
Fig. 4 shows the performance of Openia when using token representations from different layers of three models, MagicCoder 7B, CodeLlama 7B, and DeepSeek Coder 6.7B. In all three models, the first layer corresponds to layer 1, the middle layer is layer 16, and the last layer is layer 32. For each layer, we extract and evaluate the performance of Openia with the representations of tokens at four specific positions: first token and last token, which represent the starting and ending points of the entire answer generated by the LLM, and first code token and last code token, which correspond to the starting and ending points of the code segment within the LLM’s response.
In general, the deeper layer, the more stable Openia’s performance is across different token representations; conversely, the shallower layer, the more sensitive Openia becomes to variations in token selection. Specifically, when using the representations from the last layers, Openia consistently achieves the highest F1-Score with the representations of the last code token. Furthermore, its performance across different token representations from the last layers remains relatively stable. In contrast, when using representations from the first layer, there is a noticeable performance gap between different tokens. For example, in assessing the correctness of code generated by MagicCoder for independent-unit tasks, (Fig. 4(a)), Openia achieves an F1-Score of 0.7 with the last token representations from the first layer, which is 30% higher than its performance with the first code token representations.
Indeed, the first layer operates directly on raw token embeddings, and the interactions/relationships between tokens have not yet been captured. Due to the lack of contextualization, different tokens could convey varying levels of information, leading to different impacts on Openia’s performance. This explains why Openia is highly sensitive to the token selection when using representations from the first layer of the models. Meanwhile, by the time tokens reach the last layer, their representations have been enriched with contextualized information, incorporating the relationships between tokens across the sequence. As a result, representations from the last layer are more robust and less dependent on the specific token chosen, leading to stable performance of Openia across different tokens.
However, the impact of token selection on Openia’s performance varies depending on the model and layer. For instance, MagicCoder exhibits less sensitivity to token selection, while CodeLlama is highly affected. For example, with the representations from the last layer of MagicCoder, Openia’s F1-Score in predicting the code correctness for independent-unit code generation ranges from 0.74 to 0.78. However, these figures for CodeLlama vary more widely, from 0.66 to 0.75. This highlights the importance of carefully selecting suitable tokens to extract representations to obtain the best performance of Openia. Similar to layer selection, token selection should be dynamically adapted to align with the specific architecture and characteristics of the model.
5.3 Sensitivity Analysis
5.3.1 Programming language analysis
TABLE V: Correctness assessment performance in F1-Score of the approaches across different languages
Targeted language CodeBERT CodeT5+ Openia
CPP 0.65 0.67 0.82
CS 0.74 0.75 0.80
JAVA 0.59 0.64 0.68
JS 0.66 0.80 0.86
PHP 0.54 0.65 0.71
PYTHON 0.37 0.50 0.67
SH 0.58 0.58 0.64
TS 0.72 0.82 0.88
MIX 0.66 0.67 0.72
AVERAGE 0.61 0.68 0.75
As seen in Table V, Openia consistently outperforms the traditional classification-based approaches that encode the final generated code of the LLMs using pre-trained language models such as CodeBERT and CodeT5+. Across all the studied programming languages, Openia achieves an average F1-Score of 0.75, which is 12% higher than CodeT5+ and 23% higher than CodeBERT. Notably, for Python, Openia surpasses CodeT5+ and CodeBERT by 35% and 81%, respectively. These results indicate that the internal representations of the LLMs can effectively capture not only the code semantics but also the fundamental programming principles encoded in the code generation process of the model. Levering the internal representations of the LLMs enables Openia to generalize its learned knowledge and better detect code correctness across different languages.
All the approaches achieve their highest performance when the target language of the test set is TypeScript (TS) and their lowest performance when the target language is Python. For example, Openia’s F1-Score for predicting the correctness of Python code is 0.67, which is 12% lower than its average score. Meanwhile, for TS code, Openia achieves an impressive F1-score of 0.88, which is 17% higher than its average score. A similar trend is observed with CodeBERT and CodeT5+. This is due to Python’s inherent characteristics, such as dynamic typing, indentation-based syntax, and pythonic idioms, which differ significantly from the other languages. These features introduce ambiguity and variability, making it harder for Openia to generalize its learned knowledge to Python code. In contrast, TS shares many common characteristics with other languages, such as supporting both static and dynamic typing, verbose and explicit code, and curly braces-based syntax. As a result, Openia can effectively generalize its learned knowledge from other languages to accurately predict the correctness of TS code.
5.3.2 Task difficulty
Figure 5: Openia’s performance on code generated for tasks of different difficulty levels
In this experiment, we investigate how the difficulty levels of the input tasks affect Openia’s performance. First, we prompt the LLM to classify each task in the HumanEval dataset into one of five difficulty levels: Very Easy, Easy, Medium, Hard, and Very Hard. The prompt used and the assigned difficulty levels for each task are available on our website [36]. Then, we analyze the performance of Openia in predicting the correctness of the generated code for tasks within each difficulty group.
As shown in Fig. 5, Openia obtains higher performance on both easy and hard tasks, while its performance declines for medium-level tasks. Specifically, for easy tasks, Openia correctly predicts the correctness of generated code with an F1-score of 0.75, which is 25% higher than its performance on medium-level tasks. These findings suggest that the internal representations of the LLM can capture its confidence about its responses. For instance, the LLM tends to be more confident when handling easy tasks and less confident with hard ones. This confidence signal can effectively help to assess code correctness, as reflected in Openia’s high performance on these tasks. However, for medium-level tasks, the confidence signals appear to be less distinct, leading to a decrease in Openia’s performance.
5.4 Time complexity
The time complexity of Openia is primarily influenced by the extraction of internal representations from the studied code LLMs, training the lightweight probing classifier, and the inference process.
Extracting the internal representations from the studied code LLMs was conducted on an NVIDIA A6000 GPU with 48GB of VRAM. This process required approximately 86 hours for the entire dataset of 24,890 code units generated by the LLMs. Meanwhile, all the other experiments were conducted on a single NVIDIA P100 GPU with 16GB of VRAM. Openia is highly efficient in terms of both training and inference. With a lightweight probing classifier consisting of two hidden layers (128 and 64 neurons), the training process takes approximately 1 minute, as there is no embedding computation overhead. In contrast, classifiers built on pre-trained models such as CodeBERT and CodeT5+ require significantly more time, taking about 83 minutes for training. The inference process of Openia requires only 0.6 milliseconds per generated code unit. This demonstrates the feasibility of Openia for real-time, in-process assessment of code correctness during generation by LLMs, offering substantial advantages in practical deployment scenarios.
5.5 Threats to Validity
The main threats to the validity of our work consist of internal, construct, and external threats.
Internal Validity. Threats to internal validity include the selection of hyperparameters for the baseline methods and Openia. To mitigate this threat, we systematically explored different hyperparameter configurations for Openia and evaluated its performance under varied conditions (as detailed in Section 5). For baseline methods, we reused implementations and settings from the original papers, such as those for CodeBERT and CodeT5+ [40, 42], ensuring fair and consistent comparisons. Additionally, potential threats could arise from the procedures used to analyze the internal signals of LLMs and correlate them with code correctness. To address this, we implemented robust and well-documented methods to extract and interpret the intermediate states of LLMs during code generation. These methods were validated on diverse benchmarks and cross-verified using established practices in the field. To further minimize the risk of errors in implementation, our codebase has undergone extensive internal review, and we have made it publicly available [36].
Construct Validity. Threats to construct validity relate to the appropriateness of our evaluation metrics and procedures. We employed widely recognized metrics, including Accuracy, Precision, Recall, and F1-Score, to measure code correctness, as these are standard benchmarks in evaluating code generation models. Additionally, we designed experiments to include both standalone and repository-level code generation tasks, ensuring comprehensive evaluations. However, our evaluation primarily relies on automated metrics and does not yet incorporate human assessments. While automated metrics provide objectivity and scalability, human evaluation could offer deeper insights into the practical utility of Openia in real-world coding scenarios. We plan to include such assessments in future work, engaging experienced developers to evaluate the generated code based on factors like functionality, readability, and adherence to programming standards. Another potential limitation is that our controlled experimental setup may not fully reflect real-world software development environments. To address this, we evaluated Openia on multiple representative benchmarks, covering diverse programming constructs and code generation requirements.
External Validity. Threats to external validity concern the generalizability of our results to different LLMs, programming languages, and application domains. To mitigate this, we selected three widely used and representative open-source code LLMs: DeepSeek-Coder, CodeLlama, and MagicCoder [16, 17, 18]. These models are recognized for their strong performance in diverse code-related tasks, making our findings broadly applicable within the code generation domain. A further potential threat arises from focusing primarily on models with parameters up to 7B due to hardware limitations. While this constraint ensures feasibility within our experimental setup, it may limit the applicability of our findings to larger models. Future work will explore scalability to larger models and extend experiments to a broader range of programming languages and codebases to enhance generalizability.
6 Related Work
LLM-based Code Generation. Code generation is an essential task in software development, aimed at assisting programmers by suggesting subsequent code units based on the current context [46, 47, 48, 49, 50]. Recently, Large Language Models for code (Code LLMs) [17, 16, 51, 52] such as Codedex, Code Llama, and Deepseek Coder have emerged as promising tools for code generation/completion, offering the potential to automate repetitive tasks and increase developer productivity. These models, trained on massive datasets and equipped with billions of parameters, have demonstrated remarkable performance in code generation and completion [53, 54, 55]. Code LLMs have been deployed as auto-completion plugins, such as Github Copilot and CodeGeeX2 [56] in modern IDEs, and successfully streamline real-world software development activities to a certain degree [57, 58, 22, 59]. Beyond standard code generation and completion, repo(sitory)-level code generation adds another layer of complexity as it requires integrating repository-specific elements such as user-defined APIs, inter-module dependencies, and project-specific code conventions [45, 58, 60, 57]. To address the challenges of repo-level code completion, recent approaches have applied the Retrieval-Augmented-Generation (RAG) strategy. By leveraging repository-specific knowledge, RAG-based methods enhance the capabilities of Code LLMs, significantly improving their performance in completing repo-level code tasks [45, 60, 61, 62, 44].
Quality Assurance for LLM-generated Code. The quality of code generated by LLMs has garnered significant attention due to the increasing reliance on AI-generated code in real-world applications. Several empirical studies have investigated bugs and security issues in LLM-generated code, highlighting the potential risks associated with their use [63, 64, 14, 65, 66]. These studies provide a detailed analysis of common flaws and vulnerabilities, emphasizing the need for robust evaluation and mitigation strategies.
Hallucinations, where LLMs produce plausible yet incorrect or non-functional code, are another prominent issue in practical code generation [67]. Understanding and addressing these hallucinations is essential for ensuring the reliability of LLMs in software development. Furthermore, research has shown that AI-generated code often contains vulnerabilities, necessitating a deeper investigation into its safety implications [68]. For instance, comparative studies have assessed the security vulnerabilities of ChatGPT-generated code against traditional sources like StackOverflow, revealing notable differences in the safety and correctness of the generated code [69]. Efforts have also been made to analyze the cognitive processes of LLMs during code generation. A study explored whether LLMs pay attention to code structures in a manner similar to human programmers, providing insights into their decision-making and error tendencies [70]. Additionally, uncertainty analysis for LLMs has been proposed to measure their confidence in generated outputs, which could aid in identifying potentially unreliable code [71].
To enhance the reliability of LLM-generated code, researchers have introduced methods like “slow-thinking,” which instructs LLMs to perform step-by-step analyses of code functionality to detect errors more effectively. Multi-agent frameworks have also been developed to secure code generation using static analysis and fuzz testing, such as AutoSafeCoder [72]. Similarly, LLMSecGuard combines static code analyzers with LLMs to provide enhanced code security, leveraging the strengths of both approaches to mitigate vulnerabilities and ensure code robustness [73].
Bug/Vulnerability Detection in Code. Detecting bugs and vulnerabilities in code is a critical aspect of software quality assurance, aiming to identify potential issues before they lead to security breaches or system failures. Various methods have been proposed to assess the vulnerability of code components, such as files, functions, methods, or even individual lines of code [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. Devign[78] employs a graph-based approach leveraging the Code Property Graph (CPG) to predict vulnerabilities effectively. Similarly, tools like VulDeePecker[76] and SySeVR [87] focus on detecting vulnerabilities at the slice level, enabling more fine-grained identification. To enhance the granularity of vulnerability detection, graph-based models have also been developed. IVDetect[88] applies a graph-based neural network to detect vulnerabilities at the function level while using interpretable models to pinpoint vulnerable statements within suspicious functions. Similarly, VelVet[89] adopts graph-based techniques to detect vulnerabilities directly at the statement level.
More recent approaches integrate pre-trained models like CodeBERT [40] and CodeT5 [41], which have demonstrated significant effectiveness in both function- and statement-level vulnerability detection tasks. For example, LineVul[90] and LineVD[91] leverage CodeBERT to encode code representations, outperforming traditional methods like IVDetect in identifying vulnerabilities at finer levels of granularity. The success of these pre-trained models is largely attributed to their ability to capture rich contextual information and encode both syntactic and semantic characteristics of code efficiently. As a result, models like CodeBERT and CodeT5 have established themselves as powerful baselines for advancing the field of bug and vulnerability detection, driving progress toward more accurate and effective software quality assurance tools.
Analyzing LLM Internal States in Hallucination Detection and Mitigation. Hallucinations, where LLMs generate plausible but inaccurate or false information, remain a critical challenge in the deployment of these models for real-world applications [19, 20]. TrustLLM [92] provides a comprehensive study on trustworthiness in LLMs, covering multiple dimensions such as trust principles, benchmarks, and evaluation methods for mainstream LLMs. Recently, several studies have introduced white-box/open-box approaches in detecting and mitigating hallucination of general-purpose LLMs, which require analyzing the internal states of LLMs to uncover the root causes and enhance their reliability [93, 25]. Azaria et al.  [94] introduce a classifier trained on the hidden layer activations of LLMs that can predict the probability of a statement being truthful. Another study by Ji et al.  [43] explores the internal mechanisms across diverse Natural Language Generation (NLG) tasks, analyzing over 700 datasets to uncover patterns in LLM behavior. INSIDE [95] leverages the dense semantic information within LLMs’ internal states to detect hallucinations. FactoScope[96] further investigates factual discernment by measuring the inner states of LLMs, providing insights into their capacity to differentiate truth from hallucination.
Beyond analyzing internal states, alternative methods have been proposed to enhance the truthfulness and reliability of LLM outputs. LLMGuardrail [97] integrates causal analysis and adversarial learning to develop unbiased steering representations for improving output quality. LookBack [98] uses attention maps to identify and mitigate contextual hallucinations. Fact-checking through token-level uncertainty quantification [99] provides a complementary approach for verifying the factual accuracy of LLM-generated outputs. Inference-Time Intervention (ITI) [100] focuses on enhancing LLM truthfulness by dynamically modifying model activations during inference.
7 Conclusion
In this work, we introduced Openia, a novel framework that leverages the internal representations of code LLMs to assess the correctness of generated code. Unlike traditional black-box approaches, Openia adopts a white-box methodology, systematically analyzing the intermediate signals of code-specific LLMs such as DeepSeek-Coder, CodeLlama, and MagicCoder. Our empirical findings highlight that these internal signals capture latent information, including key programming principles, which strongly correlate with the reliability of the generated code. Openia not only provides a more nuanced and reliable evaluation of code correctness but also demonstrates adaptability and robustness across varying benchmarks and tasks. By unlocking the potential of in-process signals, Openia bridges the gap between code generation and quality assurance, offering a proactive approach to enhancing the reliability of LLM-generated code. Our framework sets the foundation for future research into leveraging the latent capabilities of LLMs to improve AI-driven software development.